The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: Eva K. Grebel, You-Hua Chu, and Dominik J. Bomans

                                 mcnews@astro.uiuc.edu
                      http://www.astro.uiuc.edu/mcnews/MCNews.html
                         http://www.astro.uni-bonn.de/~mcnews/

No. 19 April 5, 1998

Table of Contents

News  1
Abstracts of 9 refereed papers  2
Abstracts of 6 non-refereed papers 9

News

The third announcement for IAU Symposium 190, “New Views of the Magellanic Clouds” in Victoria, Canada, is now available. The third announcement contains the scientific program and the speakers for IAU Symposium 190. The Symposium will take place in only three months from now on July 13 – 17, 1998.

Please note that April 15 is the deadline for early registration and payment, abstracts, and accommodation reservation.

We encourage you to view the Symposium web page at http://cadcwww.hia.nrc.ca/iau190/. Here you will find detailed information of the meeting venue, travel to Victoria, planned special events, and accommodation arrangements. You will also find convenient WWW versions of the abstract submission and registration forms. This site will also be updated regularly with details of the scientific program if further changes are made. If you have any questions please don’t hesitate to contact the

organizers at iau190@hia.nrc.ca.

We expect an exciting meeting with many new results and an emerging comprehensive multiwavelength view of the Magellanic Clouds. We are looking forward to seeing you in Victoria!

Abstracts of Refereed Papers

Results of the SEST Key Programme: CO in the Magellanic Clouds VII. 30 Doradus and its southern H ii regions

L.E.B. Johansson^{1}, A. Greve ^{2}, R.S. Booth^{1}, F. Boulanger^{3,4}, G. Garay^{5},
Th. de Graauw^{6}, F.P. Israel^{7}, M.L. Kutner^{8,9}, J. Lequeux^{3,10}, D.C. Murphy^{11},
L.–Å. Nyman^{12} and M. Rubio^{5}

1 Onsala Space Observatory, S-439 92 Onsala, Sweden 2 IRAM, 300 rue de la Piscine, Domaine Universitaire, F-38406 St. Martin d’Heres, France 3 Radioastronomie, Ecole Normale Superieure, 24 Rue Lhomond, F-75231 Paris CEDEX 05, France 4 Institut d’Astrophysique Spatiale, Bat. 120, Université de Paris-XI, F-91045 Orsay CEDEX, France 5 Departamento de Astronomía, Universidad de Chile, Casilla 36-D, Santiago, Chile 6 Laboratorium voor Ruimteonderzoek, SRON, Postbus 800, NL-9700 AV Groningen,The Netherlands 7 Sterrewacht, Postbus 9513, NL-2300 RA Leiden, The Netherlands 8 NRAO, 949 N. Cherry Av., Campus Building 65, Tucson, AZ 85721-0655, USA 9 Department of Physics, Rensselaer Polytechnic Institute, Troy, NY 12180, USA 10 DEMIRM, Observatoire de Paris, 61 Av. de l’Observatoire, F-75014 Paris, France 11 The Observatories of the Carnegie Institution of Washington, 813 Santa Barbara St., Pasadena, CA 91101, USA 12 European Southern Observatory, Casilla 19001, Santiago 19, Chile

We have mapped the ^{12}CO(1–0) emission from the 30 Doradus region in the Large Magellanic Cloud with the Swedish–ESO Submillimetre Telescope (SEST). The regions investigated include the central part of the 30 Dor nebula, and the southern H ii regions N 158C, N 159, and N 160. In addition, a few prominent CO clouds have been studied in the (2–1) and (3–2) transitions of CO.

The southern area shows the strongest ^{12}CO(1–0) emission, a factor of 3 higher than towards the central part of 30 Dor. A non–LTE analysis of the CO emission from a sample of clouds indicate kinetic temperatures between 10 and 50 K; the highest temperatures are found close to the 30 Dor nebula. We have estimated the CO–H_{2} conversion factor for the two areas, separately, under the assumption that the virial mass, determined from CO parameters, reflects the total molecular mass. We find an unexpectedly small difference between the two areas. This is explained as a bias effect.

Accepted by: Astronomy & Astrophysics For preprints, contact leb@oso.chalmers.se

Ionization Structure in the 30 Doradus Nebula as seen with HST/WFPC-2

P.A. Scowen^{1}, J.J. Hester^{1}, R. Sankrit^{1}, J.S. Gallagher^{2}, G.E. Ballester^{3}, C.J. Burrows^{4}, J.T. Clarke^{3}, D. Crisp^{5}, R.W. Evans^{5}, R.E. Griffiths^{6}, J.G. Hoessel^{2}, J.A. Holtzman^{7},

J. Krist^{4}, J.R. Mould^{8}, K.R. Stapelfeldt^{5}, J.T. Trauger^{5}, A.M. Watson^{9}, J.A. Westphal^{10}

1 Department of Physics & Astronomy, Box 871504, Arizona State University, Tyler Mall, Tempe, AZ 85287-1504, USA 2 Department of Astronomy, University of Wisconsin – Madison, 475 N. Charter St., Madison, WI 53706, USA 3 Department of Atmospheric and Oceanic Sciences, University of Michigan, 2455 Hayward, Ann Arbor, MI 48109, USA 4 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 5 Jet Propulsion Laboratory, 4800 Oak Grove Drive, Pasadena, CA 91109, USA 6 Dept. of Physics - Wean Hall, Carnegie Mellon University, 5000 Forbes Ave., Pittsburgh, PA 15213-3890, USA 7 Box 30001 / Dept. 4500, New Mexico State University, Las Cruces, NM 88003, USA 8 Mount Stromlo and Siding Springs Observatories, Australian National University, Weston Creek Post Office, ACT 2611, Australia 9 Instituto de Astronomía UNAM, J. J. Tablada 1006, Col. Lomas de Santa Maria, 58090 Morelia, Michoacan, Mexico 10 Division of Geological and Planetary Sciences, California Institute of Technology, Pasadena, CA 91125, USA

Using the Hubble Space Telescope and WFPC2 we have imaged the central 20pc of the giant H II region 30 Doradus nebula in three different emission lines. The images allow us to study the nebula with a physical resolution that is within a factor of two of that typical of ground based observations of Galactic Hii regions. We present a gallery of interesting objects within the region studied. These include a tube blown by the wind of a high velocity star and a discrete Hii region around an isolated B star. This small isolated Hii region appears to be in the midst of the champagne flow phase of its evolution.

Most of the emission within 30 Dor is confined to a thin zone located between the hot interior of the nebula and surrounding dense molecular material. This zone appears to be directly analogous to the photoionized photoevaporative flows that dominate emission from small, nearby Hii regions. For example, a column of material protruding from the cavity wall to the south of the main cluster is found to be a direct analog to elephant trunks in M16. Surface brightness profiles across this structure are very similar to surface brightness profiles taken at ground based resolution across the head of the largest column in M16. The dynamical effects of the photoevaporative flow can be seen as well. An arcuate feature located above this column and a similar feature surrounding a second nearby column are interpreted as shocks where the photoevaporative flow stagnates against the high temperature gas that fills the majority of the nebula. The ram pressure in the photoevaporative flow, derived from thermal pressure at the surface of the column, is found to balance with the pressure in the interior of the nebula derived from previous X-ray observations.

By analogy with the comparison of ground and HST images of M16 we infer that the same sharply stratified structure seen in HST images of M16 almost certainly underlies the observed structure in 30 Dor. 30 Doradus is a crucial case because it allows us to bridge the gap between nearby Hii regions and the giant Hii regions seen in distant galaxies. The real significance of this result is that it demonstrates that the physical understanding gained from detailed study of photoevaporative interfaces in nearby Hii regions can be applied directly to interpretation of giant Hii regions. Stated another way, interpretation of observations of giant Hii regions must account for the fact that this emission arises not from expansive volumes of ionized gas, but instead from highly localized and extremely sharply stratified physical structures.

Accepted by: The Astronomical Journal, to appear July 1998 For preprints, contact paul.scowen@asu.edu

Also available from the URL http://tycho.la.asu.edu/html/scowen/30dor/30dor web.html
or by anonymous ftp at ftp://tycho.la.asu.edu/pub/scowen/30dor

Hubble Space Telescope Observations of Planetary Nebulae in the Magellanic Clouds VII: Cycle 3 and HST Archive Narrowband [O III] 500.7 nm Imaging

E. Vassiliadis^{1}, M.A. Dopita^{2}, S.J. Meatheringham^{3}, R.C. Bohlin^{4}, H.C. Ford^{5}, J.P. Harrington^{6}, P.R. Wood^{2}, T.P. Stecher^{7}, S.P. Maran^{8}

1Instituto de Astrofísica de Canarias, C/- Via Lactea S/N, E-38200, La Laguna, Tenerife, Spain 2Mount Stromlo and Siding Spring Observatories, Institute of Advanced Studies, The Australian National University, Private Bag, Weston Creek P.O., Weston Creek, ACT 2611, Australia 3Information Technology Services Centre, University College, Australian Defense Force Academy, Canberra, ACT 2600, Australia 4Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 5Johns Hopkins University, Department of Physics & Astronomy, Charles & 34th St, Baltimore, MD 21218, USA 6University of Maryland, Department of Astronomy, College Park, MD 20742-2421, USA 7NASA Goddard Space Flight Center, Code 680, Greenbelt, MD 20771, USA 8NASA Goddard Space Flight Center, Code 600, Greenbelt, MD 20771, USA

We present the continuation of our analysis of narrowband [O III] images, obtained using the Hubble Space Telescope, of planetary nebulae (PNs) in the Magellanic Clouds. Six objects, five in the Large Magellanic Cloud, are observed as part of the General Observer program. An additional fifteen objects are retrieved from the Hubble Space Telescope archive. The majority of the images are obtained prior to the first maintenance and refurbishment mission and are dominated by nebular [O III] 500.7 nm emission: the PN central stars are generally not seen. The raw images are deconvolved using 100 iterations of the Richardson-Lucy image restoration algorithm. The nebular dimensions and morphologies are determined. The nebular dynamical ages are derived using [O III] 500.7 nm expansion velocities from the literature. Using the techniques presented in Paper IV, the measured nebular ages are compared with theoretical evolutionary ages for H-burning and He-burning PN nuclei. Approximately 50% of the PN central stars are He-burners, in both the Large and Small Magellanic Clouds.

Accepted by: The Astrophysical Journal (August 1998) For preprints, contact vass@mso.anu.edu.au

Discovery of an Ultra-fast X-ray Pulsar in the Supernova Remnant N157B

F.E. Marshall^{1}, E.V. Gotthelf^{1}, W. Zhang^{1}, J. Middleditch^{2}, & Q.D. Wang^{3}

1 NASA / Goddard Space Flight Center, Code 660, Greenbelt, MD 20771, USA 2 Los Alamos National Lab, MS B265, CIC-19, Los Alamos, NM 87545, USA 3 Dearborn Observatory, Northwestern University, 2131 Sheridan Rd., Evanston, IL 60208, USA

We present the serendipitous discovery of 16 ms pulsed X-ray emission from the Crab-like supernova remnant N157B in the Large Magellanic Cloud. This is the fastest spinning pulsar associated with a supernova remnant (SNR). Observations with the Rossi X-ray Timing Explorer (RXTE), centered on the field containing SN1987A, reveal an X-ray pulsar with a narrow pulse profile. Archival ASCA X-ray data confirm this detection and locate the pulsar within 1′ of the supernova remnant N157B, 14′ from SN1987A. The pulsar manifests evidence for glitch(es) between the RXTE and ASCA observations which span 3.5 years; the mean linear spin-down rate is Ṗ = 5.126×10^{−14} s/s. The background subtracted pulsed emission is similar to other Crab-like pulsars with a power law of photon index of ∼ 1.6. The characteristic spin-down age (∼ 5000 years) is consistent with the previous age estimate of the SNR. The inferred B-field for a rotationally powered pulsar is ∼ 1 × 10^{12} Gauss. Our result confirms the Crab-like nature of N157B the pulsar is likely associated with a compact X-ray source revealed by ROSAT HRI observations.

Submitted to: The Astrophysical Journal Letters

For preprints, contact gotthelf@gsfc.nasa.gov
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9803214

Ultraviolet Imaging Telescope Observations of the Magellanic Clouds

Joel Wm. Parker^{1}, Jesse K. Hill^{2}, Robert Cornett^{2}, Joan Hollis^{2}, Emily Zamkoff^{2}, Ralph C. Bohlin^{3}, Robert W. O’Connell^{4}, Susan G. Neff^{5}, Morton S. Roberts^{6}, Andrew M. Smith^{5}, and Theodore P. Stecher^{5}

1Southwest Research Institute, Suite 426, 1050 Walnut Street, Boulder, CO 80302, USA 2Hughes STX Corporation, Code 681, GSFC/NASA, Greenbelt, MD 20771, USA 3Space Telescope Science Institute, Baltimore, MD 21218, USA 4Astronomy Department, University of Virginia, P.O. Box 3818, Charlottesville, VA 22903, USA 5Laboratory for Astronomy and Solar Physics, Code 681, GSFC/NASA, Greenbelt, MD 20771, USA 6National Radio Astronomy Observatory, Edgemont Road, Charlottesville, VA 22903, USA

We present wide-field far-ultraviolet (FUV: 1300–1800 Å) images of the Large and Small Magellanic Clouds (LMC and SMC). These data were obtained by the Ultraviolet Imaging Telescope (UIT ) during the Astro-1 (1990 December 1–10) and Astro-2 (1995 March 2–18) missions; the images provide an extensive FUV mosaic of the SMC (Cornett et al. 1997) and contain numerous regions in the LMC, covering a wide range of stellar densities and current star formation activity. A total of 47 LMC/Lucke-Hodge (1970) and 37 SMC/Hodge (1985) OB associations are completely or partially included in the observed fields. FUV data can identify the hottest OB stars more easily than optical photometry can, and these stars dominate the ionizing flux, which is correlated to the observed Hα flux of the associated Hii regions. Of the Hii regions in the DEM catalog (Davies, Elliott, & Meaburn 1976), the UIT fields completely or partially include 102 DEM regions in the LMC and 74 DEM regions in the SMC.

We present a catalog of FUV magnitudes derived from point spread function photometry for 37,333 stars in the LMC (the UIT FUV magnitudes for 11,306 stars in the SMC were presented in the paper by Cornett et al. 1997), with a completeness limit of m_{UV} ≈ 15 mag and a detection limit of m_{UV} ≈ 17.5. The average uncertainty in the photometry is ∼ 0.1 mag. The full catalog with astrometric positions, photometry, and other information is available via the electronic journal as well as other publicly available astronomical data archives.

We divided the catalog into field stars and stars that are in DEM regions. We analyzed each of these two sets of stars independently, comparing the composite UV luminosity function of our data to UV magnitudes derived from stellar evolution and atmosphere models in order to derive the underlying stellar formation parameters. We find a most probable initial mass function slope for the LMC field stars of Γ = −1.80 ± 0.09. The statistical significance of this single slope for the LMC field stars is extremely high, though we also find some evidence for a field star initial mass function (IMF) slope of Γ ∼ −1.4, equal to the Salpeter (1955) slope. However, in the case of the stars in the DEM regions (the stars in all the regions were analyzed together as a single group), we find three IMF slopes of roughly equal likelihood: Γ = −1.0, −1.6, and −2.0.

No typical age for the field stars is found in our data for time periods up to a continuous star formation age of 500 Myr, which is the maximum age consistent with the completeness limit magnitude of the catalog’s luminosity function. The best age for the collection of cluster stars was found to be t_{0} = 3.4 ± 1.9 Myr; this is consistent with the age expected for a collection of OB stars from many different clusters.

Accepted by: The Astronomical Journal

For preprints, contact joel@boulder.swri.edu
Also available from the URL http://www.boulder.swri.edu/~joel/papers.html
or by anonymous ftp at ftp.boulder.swri.edu/pub/joel/uit mc.ps.gz

Ages and Metallicities of Star Clusters and Surrounding Fields in the Outer Disk of the Large Magellanic Cloud

Eduardo Bica^{1}, Doug Geisler^{2}, Horacio Dottori^{1}, Juan J. Clariá^{3}, Andres E. Piatti^{3}, & João F.C. Santos Jr.^{4}

1 Departamento de Astronomia, Instituto de Física, UFRGS, C.P. 15051, 91501-970 Porto Alegre RS, Brazil 2 Kitt Peak National Observatory, National Optical Astronomy Observatories, P.O. Box 26732, Tucson, Arizona 85726, USA 3 Observatorio Astronómico de Córdoba, Laprida 854, 5000, Córdoba, Argentina 4 Dep. de Fisica, ICEx, UFMG, C.P. 702, 30123-970 Belo Horizonte MG, Brazil

We present Washington system C, T_{1} color-magnitude diagrams of 13 star clusters and their surrounding fields which lie in the outer parts of the LMC disk (r > 4^{∘}), as well as a comparison inner cluster. The total area covered is large (2/3^{□^{∘}}), allowing us to study the clusters and their fields individually and in the context of the entire galaxy. Ages are determined by means of the magnitude difference δT_{1} between the giant branch clump and the turnoff, while metallicities are derived from the location of the giant and subgiant branches as compared to fiducial star clusters. This yields a unique dataset in which ages and metallicities for both a significant sample of clusters and their fields are determined homogeneously. We find that in most cases the stellar population of each star cluster is quite similar to that of the field where it is embedded, thus sharing its mean age and metallicity. The old population (t≥10 Gyr) is detected in most fields as a small concentration of stars on the horizontal branch blueward and faintward of the prominent clump.

Three particular fields present remarkable properties: (i) The so far unique cluster ESO121-SC03 at ≈9 Gyr has a surrounding field which shares the same properties, which in turn is also unique in the sense that such a dominant old field component is not present elsewhere, at least not significantly in the fields as yet studied. (ii) The field surrounding the far eastern intermediate age cluster OHSC37 is noteworthy in the sense that we do not detect any evidence of LMC stars: it is essentially a Galactic foreground field. We can thus detect the LMC field out to > 11^{∘} (the deprojected distance of ESO121SC03), or ∼ 11 kpc, but not to 13^{∘}(∼ 13 kpc), despite the presence of clusters at this distance. (iii) In the northern part of the LMC disk the fields of SL388 and SL509 present colormagnitude diagrams with a secondary clump ≈0.45 mag fainter than the dominant intermediate age clump, suggesting a stellar population component located behind the LMC disk at a distance comparable to that of the SMC. Possibly we are witnessing a depth effect in the LMC, and the size of the corresponding structure is comparable to the size of a dwarf galaxy. The unusual spatial location of the cluster OHSC37 and the anomalous properties of the SL388 and SL509 fields might be explained as debris from previous LMC interactions with the Galaxy and/or the SMC.

The mean metallicity derived for the intermediate age outer disk clusters is <[Fe/H]>= −0.7 and for their surrounding fields <[Fe/H]>= −0.6. These values are significantly lower than found by Olszewski et al. (1991, AJ, 101, 515) for a sample of clusters of similar age, but are in good agreement with several recent studies. A few clusters stand out in the age–metallicity relation in the sense that they are intermediate age clusters at relatively low metallicity ([Fe/H]≈ −1).

Accepted by: The Astronomical Journal (July issue) For preprints, contact dgeisler@noao.edu

Also available on astro-ph/9803167

Clustering Properties of Stars in Simulations of Wind-Driven Star Formation

John Scalo^{1} and David Chappell^{1}

1 University of Texas, Austin, TX 78712, USA

Several recent observational studies have shown that the clustering of young stars in local starforming regions, and of Cepheids in the LMC, can be described by a power law two-point correlation function. We show by numerical simulations that the observed range in power law slopes can be accounted for by a model in which stellar winds drive expanding shells that are subjected to nonlinear fluid advection and interactions with other shells, and in which star formation occurs when a threshold shell column density is exceeded. The models predict how the power law slope should depend on the maximum age of the stellar sample and the average star formation rate, although a number of effects preclude a comparison with currently-available data. Stellar migration flattens the power law slope below a scale that depends on the velocity dispersion and age of the sample, an effect which may explain the secondary breaks in the observed correlation functions of some regions at large separations.

Submitted to: Astrophysical Journal (Letters) For preprints, contact John Scalo parrot@astro.as.utexas.edu

or by anonymous ftp at: ftp astro.as.utexas.edu, /us2/spool/ftppublic/pub/betty

A “Short” Distance to the Large Magellanic Cloud With the Hipparcos Calibrated Red Clump Stars

K.Z. Stanek^{1}, D. Zaritsky^{2} and J. Harris^{2}

1 Harvard-Smithsonian Center for Astrophysics, 60 Garden St., MS20, Cambridge, MA 02138, USA 2 UCO/Lick Observatory and Department of Astronomy and Astrophysics, University of California, Santa Cruz, CA 95064, USA

Following an approach developed by Paczyński & Stanek, we derive a distance to the Large Magellanic Cloud (LMC) by comparing red clump stars from the Hipparcos catalog with the red clump stars observed in two fields in the LMC that were selected from the ongoing photometric survey of the Magellanic Clouds to lie in low extinction regions. The use of red clump stars allows a single step determination of the distance modulus to the LMC, μ_{0,LMC} = 18.065±0.031±0.09 mag (statistical plus systematic error), and the corresponding distance, R_{LMC} = 41.02±0.59±1.74 kpc. This measurement is in excellent agreement with the recent determination by Udalski et al., also based on the red clump stars, but is ∼ 0.4 mag smaller than the generally accepted value of μ_{0,LMC} = 18.50 ± 0.15 mag. We discuss possible reasons for this discrepancy and how it can be resolved.

Submitted to: Astrophysical Journal Letters

For preprints, contact kstanek@cfa.harvard.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9803181

Cepheid Distance to M100 in Virgo Cluster

D. Narasimha^{1} and Anwesh Mazumdar^{1}

1 Department of Astronomy and Astrophysics, Tata Institute of Fundamental Research, Homi Bhabha Road, Mumbai 400005, India

A measurement of the distance to Virgo Cluster by a direct method along with a realistic error analysis is important for a reliable determination of the value of Hubble Constant. Cepheid variables in the face-on spiral M100 in the Virgo Cluster were observed with the Hubble Space Telescope in 1994 under the HST Key Project on the Extragalactic Distance Scale. This work is a reanalysis of the HST data following our study of the Galactic Cepheids (in an accompanying communication: “Some Characteristics of Galactic Cepheids Relevant to the Calibration of the Distance Scale”, submitted to ApJ, also available at http://xxx.lanl.gov/abs/astro-ph/9803194). The periods of the Cepheids are determined using two independent methods and the reasons for varying estimates are analyzed. The log(period) vs V -magnitude relation is re-calibrated using LMC data as well as available HST observations for three galaxies and the slope is found to be −3.45 ± 0.15. A prescription to compute correction for the flux-limited incompleteness of the sample is given and a correction of 0 to 0.28 magnitude in V -magnitude for Cepheids in the period range of 35 to 45 days is applied. The extinction correction is carried out using period vs mean (V − I) color and V -amplitude vs (V − I) color at the brightest phase relations. The distance to M100 is estimated to be 20.4 ± 1.7 (random) ±2.4 (systematic) Mpc.

Submitted to: The Astrophysical Journal

For preprints, contact anwesh@astro.tifr.res.in
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9803195

Abstracts of Non-Refereed Papers

The ATCA Radio-continuum Investigation of the Magellanic Clouds

Miroslav D. Filipović^{1,2,3} and Lister Staveley-Smith^{3}

1 Max-Planck-Institut für extraterrestrische Physik, D–85740 Garching bei München, Germany 2 University of Western Sydney Nepean, P.O. Box 10, Kingswood, NSW 2747, Australia 3 Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW 2121, Australia

We present preliminary results from the following recent investigations of radio-continuum emission in the Magellanic Clouds: (1) a pilot survey of ∼80 supernova remnants (SNRs) and SNR candidates carried out using the Australia Telescope Compact Array (ATCA); (2) an ATCA mosaic of the Small Magellanic Cloud (SMC) at λ=13 cm.

To appear in: Proceedings of the workshop ”The Magellanic Clouds and Other Dwarf Galaxies”, Bad Honnef, January 18-22, 1998; eds. T. Richtler & J. Braun For preprints, contact fica@mpe.mpg.de

DENIS Observations on the Magellanic Clouds

M.R. Cioni^{1}, C. Loup^{2} H. Habing^{1}, and E. Deul^{1}

1 Leiden Observatory, Niels Bohrweg 2, 2333 CA Leiden, The Netherlands 2 Institut d’Astrophysique de Paris, 98 bis Boulevard Arago, F-75014 Paris, France

The Deep Near Infrared Southern Sky Survey (DENIS) instrument (Epchtein et al., 1997) is surveying the southern hemisphere using the ESO 1m telescope in the I, J and K photometric bands and with limiting magnitudes of 18 mag in I (0.8 μm), 16 mag in J (1.25 μm) and 14 mag in K_{s} (2.15 μm). The survey is made strip by strip where each strip covers a region of 30^{o} × 12′. Large parts of both Magellanic Clouds have already been covered. We expect to find of the order of 10^{7} point sources. A first inspection shows already considerable differences in AGB population between the Large and the Small Cloud and with the bulge of our Milky Way. From a comparison of our color-color diagram and the one in the Galaxy (Schultheis et al., 1997) we conclude that there is a greater population of M-type stars in the Galaxy while in the Magellanic Clouds C-type stars seem to dominate. We believe that with all the reduced strips, that will completely cover! the Clouds (about 90 for the LMC and 50 for the SMC), this behaviour will strongly emerge. Each object as part of a final catalogue will constitute the more complete catalogue of point sources in the Magellanic Clouds at these frequencies.

To appear in: Proceeding of the workshop “The Magellanic Clouds and Other Dwarf Galaxies”, Bad Honnef, Jan 18-22, 1998; eds. T. Richtler & J. Braun

For preprints, contact mrcioni@strw.leidenuniv.nl

The MACHO Project LMC Variable Star Inventory: Classical Cepheids and AGB Variables

D. Alves^{1}, C. Alcock^{1}, K. Cook^{1}, S. Marshall^{1}, D. Minniti^{1}, R. Allsman^{2}, T. Axelrod^{2},

K. Freeman^{2}, B. Peterson^{2}, A. Rodgers^{2}, K. Griest^{3}, T. Vandehei^{3}, A. Becker^{4}, C. Stubbs^{4}, A. Tomaney^{4}, D. Bennett^{5}, M. Lehner^{6}, P. Quinn^{7}, M. Pratt^{8},

W. Sutherland^{9}, D. Welch^{10}

1 Lawrence Livermore National Laboratory, Livermore, CA 94550, USA 2 Mt. Stromlo and Siding Spring Observatories, Australian National University, Weston Creek, ACT 2611, Australia 3 Department of Physics, University of California, San Diego, La Jolla, CA 92093, USA 4 Departments of Astronomy and Physics, University of Washington, Seattle, WA 98195, USA 5 Physics Department, University of Notre Dame, Notre Dame, IN 46556, USA 6 University of Sheffield, UK 7 European Southern Observatory, Karl-Schwarzschild Str. 2, D-85748, Garching, Germany 8 MIT, USA 9 Department of Physics, University of Oxford, Oxford OX1 3RH, U.K. 10 Dept. of Physics & Astronomy, McMaster University, Hamilton, Ontario, L8S 4M1 Canada

The MACHO Project’s time-sampled photometric database contains thousands of LMC variable star lightcurves. We briefly review the Cepheid catalog. A third sequence is identified in the period– luminosity diagram, lying parallel to and fainter than the 1H and F mode sequences. These are excellent candidates to be Cepheids crossing the instability strip for the first time. However, we find no evidence for period-changes in the four-year MACHO lightcurves of these Cepheids. Anticipating the MACHO four-year LMC variable star catalog, we undertake a preliminary reconnaissance of the ∼18,500 year-one “red variables”. The vast majority are newly discovered low-amplitude variable stars, and may be on the asymptotic giant branch (AGB). We present the dual peaked differential luminosity function, period–frequency histograms, and the period–luminosity diagram. New AGB variables in the LMC clusters NGC 1783 and NGC 1898 are identified. We make a comparison with AGB variables in 47 Tuc and theoretical data; implications for the Mira PL relation are discussed. Lastly, we present the lightcurve of a carbon-rich AGB variable found near the LMC cluster NGC 2056 whose episodic fading is consistent with discrete mass-loss events.

To appear in: Invited paper presented at Joint Discussion No.24, Pulsating Stars - Recent Developments in Theory and Observation, XXIII IAU General Assembly. Eds. M. Takeuti and D.Sasselov, Universal ress For preprints, contact alves@igpp.llnl.gov Also available from the URL astro-ph/9804003

HST Study of the Stellar Populations around SN 1987A

Martino Romaniello^{1,2}, Nino Panagia^{1,3}, Salvo Scuderi^{4} and the SINS collaboration

1Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2Scuola Normale Superiore, Pisa, Italy 3On assignment from Astrophysics Division, Space Science Department of ESA 4Osservatorio Astrofisico di Catania, Viale A. Doria 6, I-95125, Catania, Italy

We present a short account of a study of the stellar population around SN 1987A based on an analysis of multi-band HST-WFPC2 images.

The effective temperature, radius and reddening of each star were determined by fitting the measured broad band magnitudes to the ones calculated with model atmospheres. In addition, all stars with Hα equivalent widths in excess of 8Å were identified.

An inspection to the HR diagram reveals the presence of several generations of stars, with ages between 1 and 150 Myrs, superposed on a much older field population. The youngest stars in the field appear to be T Tau stars, characterized by strong Hα excesses. We conclude that SN 1987A is associated with a region in which star formation has been active over a long time interval and is still very active at present.

To appear in: Proceedings of the workshop “The Large Magellanic Cloud and Other Dwarf Galaxies”, Bad Honnef, January 18-22, 1998, eds. T. Richtler & J.M. Braun For preprints, contact martino@stsci.edu

Axisymmetrical Structures of Planetary Nebulae and SN 1987A

Noam Soker

Dept. of Physics, Univ. of Haifa, Israel

I summarize some recent models and ideas for the formation of axisymmetrical structures of planetary nebulae and the three rings of SN 1987A, as follows. (a) I review the general role of binary companions, including brown dwarfs and planets. (b) I propose a mechanism for axisymmetrical mass loss on the AGB that may account for the axially symmetric structures of elliptical planetary nebulae and that operates for slowly rotating AGB stars. (c) I propose a model for the formation of the two outer rings of SN 1987A, which is based on the numerical simulation of Soker (1989), and discuss a mechanism for their displacement from the exploding star.

In the proceedings of the conference “Physical Processes in Astrophysical Fluids”. Will be published as a special volume of Physics Reports.

For preprints, contact soker@physics.technion.ac.il

MACHOs as Brown Dwarfs

Francesco De Paolis^{1}, Gabriele Ingrosso^{2}, Philippe Jetzer^{3}, and Marco Roncadelli^{4}

1 Bartol Research Institute, University of Delaware, USA 2 Dipartimento di Fisica, Università di Lecce and INFN, Italy 3 Paul Scherrer Institut, Laboratory for Astrophysics and Institute of Theoretical Physics, University of Zurich, Switzerland 4 INFN, Sezione di Pavia, Italy

Recent observations of microlensing events in the Large Magellanic Cloud suggest that a sizable fraction of the galactic halo is in the form of Massive Astrophysical Compact Halo Objects (MACHOs). Although the average MACHO mass is presently poorly known, the value ∼ 0.1M_{⊙} looks as a realistic estimate, thereby implying that brown dwarfs are a viable and natural candidate for MACHOs. We describe a scenario in which dark clusters of MACHOs and cold molecular clouds (mainly of H_{2}) naturally form in the halo at galactocentric distances larger than 10-20 kpc. Moreover, we discuss various experimental tests of this picture.

To appear in the proceedings of the workshop “DM-ITALIA-97”, Trieste, Italy

For preprints, contact jetzer@physik.unizh.ch
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9803200